# Low-Altitude ISAC with Rotatable Active and Passive Arrays

### v1

@@ Abstract
This paper investigates a low-altitude integrated sensing and communication (ISAC) system that leverages cooperative rotatable active and passive arrays. We consider a downlink scenario where a base station (BS) with an active rotatable array serves multiple communication users and senses low-altitude targets, assisted by a rotatable reconfigurable intelligent surface (RIS). A rotation-aware geometry-based multipath model is developed to capture the impact of three-dimensional (3D) array orientations on both steering vectors and direction-dependent element gains. On this basis, we formulate a new optimization problem that maximizes the downlink sum rate subject to a transmit power budget, RIS unit-modulus constraints, mechanical rotation limits, and a sensing beampattern mean-squared-error (MSE) constraint. To address the resulting highly non-convex problem, we propose a penalty-based alternating-optimization (AO) framework that alternately updates the BS precoder, RIS phase shifts, and BS/RIS array rotation angles. The three blocks are efficiently handled via a convex optimization method based on quadratic-transform (QT) and majorization-minorization (MM), Riemannian conjugate gradient (RCG) on the unit-modulus manifold, and projected gradient descent (PGD) with Barzilai–Borwein step sizes, respectively. Numerical results in low-altitude geometries demonstrate that the proposed jointly rotatable BS–RIS architecture achieves significant sum-rate gains over fixed or partially rotatable baselines while guaranteeing sensing requirements, especially with directional antennas and in interference-limited regimes.
@@ I Introduction
Low-altitude wireless networks (LAWNs) are emerging as a key component of next-generation intelligent infrastructures, interconnecting uncrewed aerial vehicles (UAVs), low-altitude platform stations, aerial robots, and connected ground terminals in the three-dimensional (3D) airspace below a few kilometers [1]-[4]. Typical LAWN applications, such as aerial surveillance, disaster response, smart-city sensing, and low-altitude traffic management, rely not only on reliable broadband connectivity, but also on accurate situational awareness of dynamic low-altitude objects. This naturally calls for integrated sensing and communication (ISAC) designs that can jointly support data transmission and high-resolution sensing within complex, rapidly changing airspace environments [5],[6]. Compared with conventional terrestrial cellular systems, LAWNs must cope with fast-varying spatial geometries and require shaping radio signals to illuminate aerial targets with stringent size–weight–power constraints and capture their reflections, which make efficient ISAC-oriented signal processing particularly challenging.
@@
Multiple-input multiple-output (MIMO) technology and advanced precoding have long been recognized as key tools to exploit spatial degrees of freedom in wireless systems [7],[8]. MIMO systems provide significant gains in beamforming, spatial multiplexing, and diversity by leveraging independent or quasi-independent channel fading, thereby significantly improving spectral efficiency and link reliability. However, in classical designs, antenna arrays are fixed in position and orientation, limiting spatial control to electronic beamforming. For LAWNs, where users and targets are distributed in full 3D and may move at low altitude with complex trajectories, purely electronic beamforming is often not sufficient: system performance becomes strongly dependent on the relative 3D geometry among the base station (BS), scatterers, and low-altitude objects. This motivates the exploitation of geometry-aware architectures that can reshape both the propagation environment and the array response itself for LAWN-specific ISAC.
@@
Reconfigurable intelligent surfaces (RISs) offer a promising approach to tailor the wireless environment via nearly passive phase-shift elements that reflect incident signals toward desired directions [9]-[11]. RIS-aided systems have been extensively investigated for coverage enhancement, interference mitigation, and energy-efficient MIMO operation. More recently, in addition to communication, RIS has been shown to improve wireless sensing, including target detection and localization, by redirecting illumination toward non-line-of-sight (NLoS) targets or acting as a passive reflector from favorable angles, thereby enriching spatial diversity [12]-[14]. In parallel, movable antennas (MA) has recently been proposed to further enhance spatial flexibility by allowing antenna elements or arrays to change their positions within a small region [15]-[18], enabling substantial gains in diversity, multiplexing, and interference management [19], and have been extended to multicast scenarios [20], as well as to two-timescale designs exploiting statistical channel state information (CSI) [21]. These works demonstrate that adding mechanical degrees of freedom to electronic beamforming can significantly improve performance.
@@
However, most existing MA- or RIS-related works treat antenna movement or passive phase control in isolation [22], and their ISAC designs for LAWN-like geometries are much less explored: existing ISAC designs typically assume fixed arrays and do not exploit low-altitude 3D geometry control to shape the transmit beampattern and the multiuser downlink channels simultaneously [23]. Furthermore, the mechanical flexibility in the MA literature is typically modeled as small-scale position changes within a local region. In contrast, realistic low-altitude platforms (e.g., UAV-mounted arrays or rooftop BS panels) can be designed to support rotation of the entire active or passive array via low-cost mechanical actuators [24]. Such rotatable arrays (RA) can jointly adjust their boresight directions, element coordinates, and directional gains, offering a distinct and practically relevant form of geometry control [25]. In summary, RIS-aided ISAC designs that consider rotatable passive arrays and their coupling with active-array rotations at the BS have yet to be systematically studied, especially under explicit constraints on the similarity of sensing beam patterns. As a result, the fundamental question of how to jointly design rotatable active and passive arrays, RIS phase shifts, and multiuser transmit beamforming for low-altitude ISAC remains open.
@@
Motivated by these observations, this paper investigates a low-altitude ISAC architecture with cooperatively rotatable active and passive arrays. We consider a downlink LAWN scenario where a BS equipped with an active rotatable array simultaneously serves multiple single-antenna ground users and senses low-altitude targets in monostatic mode, assisted by a rotatable RIS termed passive array. Both arrays can rotate in 3D, and the resulting orientations jointly affect (i) the element coordinates in the steering vectors and (ii) direction-dependent element/array gains through a practical directivity model. Based on the rotation-aware channel model, we explicitly couple communication and sensing performance and formulate a joint design problem that maximizes the sum rate subject to a sensing constraint, a BS transmit-power budget, and mechanical rotation limits at both the BS and RIS, with design variables given by the BS ISAC precoder, the RIS phase-shift vector, and the 3D rotation angles of the two arrays.
The main contributions of this work are summarized as
@@
  • We propose a LAWN-oriented ISAC architecture in which a rotatable active BS array cooperates with a rotatable passive RIS array to serve multiple communication users and jointly sense low-altitude targets. We develop a rotation-aware 3D multipath channel model that captures both phase shifts from rotated antenna/element positions and direction-dependent antenna/element directional radiation pattern gains. Based on this model, we derive closed-form expressions for the multiuser downlink SINR and the monostatic sensing beampattern, and formulate a new optimization problem that maximizes the communication sum rate under a sensing beampattern mean-squared-error (MSE) constraint, power budget on the BS, unit-modulus constraint on the RIS and mechanical rotation bounds on RA, thereby revealing mechanical rotations as additional spatial degrees of freedom for LAWN ISAC.
@@
  • We convert the constrained ISAC design into a penalized formulation with a hinge penalty on the sensing MSE and develop an alternating optimization (AO) framework that iteratively updates the BS precoder, RIS phase shifts, and array rotations. For the BS precoding block, we combine the quadratic transform (QT) and the Lagrangian dual transform (LDT) with a majorization–minimization (MM) step to obtain a tractable quadratic-constrained quadratic program (QCQP) subproblem that admits a closed-form water-filling-like update. For the RIS block, we exploit the complex-circle manifold structure and design an efficient Riemannian conjugate gradient (RCG) algorithm. For the rotation block, we derive closed-form gradients of both SINR and beampattern MSE with respect to the BS/RIS Euler angles and employ a projected gradient descent (PGD) method with Barzilai–Borwein initialization and Armijo backtracking.
@@
  • We conduct extensive simulations using LAWN-inspired 3D geometries and compare the proposed design with several baselines, including only BS rotation, only RIS rotation, fully fixed arrays, and non-RIS configurations. The results show that the jointly rotatable BS–RIS architecture provides substantial sum-rate gains under the same sensing MSE constraint across varying transmit powers, antenna array sizes, and user densities. The gains are further amplified with realistic directional element patterns, highlighting the importance of aligning the array boresights and main lobes with both user and target directions in a low-altitude 3D space.
@@
The rest of the paper is structured as follows: Section II introduces the system model and problem formulation. Section III presents proposed algorithm. Section IV presents the numerical results, and Section V concludes the paper.
@@
Figure 1: Illustration of a low-altitude ISAC system with rotatable active and passive arrays.
@@ II System Model and Problem Formulation
We consider a downlink RA-enabled ISAC system, as shown in Fig. 1, where a base station (BS) uses an active rotatable uniform planar array (UPA) with $M$ transmit antennas to serve $K$ single-antenna users while sensing multiple targets in a monostatic mode; at the same time, a passive rotatable UPA RIS with $N$ reflecting elements assists both tasks. In addition to active transmit beamforming at the BS and passive phase shifts at the RIS, both arrays can rotate in 3D. These rotations steer the array boresights and reshape the array response to improve, or balance, communication and sensing performance.
@@ II-A Array Geometry and Rotations
We adopt a global right-handed Cartesian coordinate system (CCS) $\mathcal{F}_{0}=(O,x,y,z)$. The BS employs a co-located transmit/receive UPA with size $M=M_{x}\times M_{y}$, and the RIS is a UPA with size $N=N_{x}\times N_{y}$. Let $\overline{\boldsymbol{d}}^{\mathrm{B}}_{m}\in\mathbb{R}^{3}$ and $\overline{\boldsymbol{d}}^{\mathrm{R}}_{n}\in\mathbb{R}^{3}$ denote the unrotated local coordinates of the $m$-th BS transmit antenna and the $n$-th RIS reflecting element, respectively, and let $\boldsymbol{d}_{0}^{\mathrm{B}}$ and $\boldsymbol{d}_{0}^{\mathrm{R}}$ be their array centers in $\mathcal{F}_{0}$. Both BS and RIS employ a motorized mechanism to rotate the arrays as a whole. Based on the system’s ISAC requirements, the proper rotation angles are determined by a central control center, which is connected to the motor controllers via fiber links.
Each array rotates as a rigid body through Euler angles
$\boldsymbol{r}^{\text{B}}\triangleq\big[\zeta_{x}^{\text{B}},\zeta_{y}^{\text{B}},\zeta_{z}^{\text{B}}\big]^{T},$ (1a)
$\boldsymbol{r}^{\text{R}}\triangleq\big[\zeta_{x}^{\text{R}},\zeta_{y}^{\text{R}},\zeta_{z}^{\text{R}}\big]^{T},$ (1b)
respectively, in their local CCS. The corresponding rotation matrix generated by $\boldsymbol{r}^{\text{A}}$, where we define $\text{A}\in\{\text{B,R}\}$ for brevity, represents the rotations about their local axis with the standard intrinsic $z$-$y$-$x$ convention and is given by
$\boldsymbol{R}\left(\zeta_{x}^{\text{A}},\zeta_{y}^{\text{A}},\zeta_{z}^{\text{A}}\right)=\boldsymbol{R}_{x}\left(\zeta_{x}^{\text{A}}\right)\boldsymbol{R}_{y}\left(\zeta_{y}^{\text{A}}\right),\boldsymbol{R}_{z}\left(\zeta_{z}^{\text{A}}\right),$ (2)
@@
$\boldsymbol{R}_{x}\left(\zeta_{x}^{\text{A}}\right)=\left[\begin{matrix}1&0&0\\ 0&\cos\zeta_{x}^{\text{A}}&-\sin\zeta_{x}^{\text{A}}\\ 0&\sin\zeta_{x}^{\text{A}}&\cos\zeta_{x}^{\text{A}}\\ \end{matrix}\right],$ (3a)
$\boldsymbol{R}_{y}\left(\zeta_{y}^{\text{A}}\right)=\left[\begin{matrix}\cos\zeta_{y}^{\text{A}}&0&\sin\zeta_{y}^{\text{A}}\\ 0&1&0\\ -\sin\zeta_{y}^{\text{A}}&0&\cos\zeta_{y}^{\text{A}}\\ \end{matrix}\right],$ (3b)
$\boldsymbol{R}_{z}\left(\zeta_{z}^{\text{A}}\right)=\left[\begin{matrix}\cos\zeta_{z}^{\text{A}}&-\sin\zeta_{z}^{\text{A}}&0\\ \sin\zeta_{z}^{\text{A}}&\cos\zeta_{z}^{\text{A}}&0\\ 0&0&1\\ \end{matrix}\right].$ (3c)
The rotated antenna or element coordinates in $\mathcal{F}_{0}$ are then

### v2

@@ Abstract
This paper investigates a low-altitude integrated sensing and communication (ISAC) system that leverages cooperatively rotatable active and passive arrays. We consider a downlink scenario in which a base station (BS) equipped with a three-dimensional (3D) rotatable active array serves multiple users while sensing low-altitude targets in monostatic mode, assisted by a rotatable reconfigurable intelligent surface (RIS). To capture the impact of array orientations, we develop a rotation-aware geometry-based multipath model in which 3D rotations jointly affect the phases of the steering vector via rotated element coordinates, and the direction-dependent element and antenna radiation gains via a practical directivity model. On this basis, we formulate a joint design problem that maximizes a weighted-sum utility of the downlink sum rate and the sensing beampattern similarity, where the latter is quantified by a normalized mean-squared error (NMSE). We introduce a scaling parameter and update it in closed form once per outer alternating-optimization (AO) iteration. The resulting non-convex problem is tackled via an efficient AO framework that alternately updates the BS precoder, RIS phase shifts, and array rotation angles, using a quadratic transform and majorization-minimization-based closed-form updates, Riemannian conjugate gradient on the unit-modulus manifold, and projected gradient ascent under rotation box constraints, respectively. Numerical results demonstrate that the proposed jointly rotatable arrays at the BS and RIS achieve a favorable rate-sensing trade-off and performance gains compared with fixed or partially rotatable baselines.
@@ I Introduction
Low-altitude wireless networks (LAWNs) are emerging as a key component of next-generation intelligent infrastructures, interconnecting uncrewed aerial vehicles (UAVs), low-altitude platform stations, aerial robots, and connected ground terminals in the three-dimensional (3D) airspace below a few kilometers [1, 2, 3, 4]. Typical LAWN applications, such as aerial surveillance, disaster response, smart-city sensing, and low-altitude traffic management, rely not only on reliable broadband connectivity but also on accurate situational awareness of dynamic low-altitude objects. This naturally calls for integrated sensing and communication (ISAC) designs that can jointly support data transmission and high-resolution sensing within complex, rapidly changing airspace environments [5, 6]. Compared with conventional terrestrial cellular systems, LAWNs must cope with rapidly varying spatial geometries and require shaping radio signals to illuminate aerial targets while meeting stringent size–weight–power constraints and capturing their reflections, which make efficient ISAC-oriented signal processing particularly challenging.
@@
Multiple-input multiple-output (MIMO) technology and advanced precoding have long been recognized as key tools to exploit spatial degrees of freedom (DoF) in wireless systems [7, 8]. MIMO systems provide significant gains in beamforming, spatial multiplexing, and diversity by leveraging independent or quasi-independent channel fading, thereby significantly improving spectral efficiency and link reliability. However, in classical designs, antenna arrays are fixed in position and orientation, limiting spatial control to electronic beamforming. For LAWNs, where users and targets are distributed in full 3D and may move at low altitude with complex trajectories, purely electronic beamforming is often not sufficient, as system performance becomes strongly dependent on the relative 3D geometry among the base station (BS), scatterers, and low-altitude objects. This motivates the exploitation of geometry-aware architectures that can reshape both the propagation environment and the array response itself for LAWN-specific ISAC.
@@
Reconfigurable intelligent surfaces (RISs) offer a promising approach to tailor the wireless environment via nearly passive phase-shift elements that reflect incident signals toward desired directions [9, 10, 11, 12, 13]. RIS-aided systems have been extensively investigated for coverage enhancement, interference mitigation, and energy-efficient MIMO operation. More recently, in addition to communication, RIS has been shown to improve wireless sensing, including target detection and localization, by redirecting illumination toward non-line-of-sight (NLoS) targets or acting as a passive reflector from favorable angles, thereby enriching spatial diversity [14, 15, 16]. In parallel, movable antennas (MA) has recently been proposed to further enhance spatial flexibility by allowing antenna elements or arrays to change their positions within a small region [17, 18, 19, 20], enabling substantial gains in diversity, multiplexing, and interference management [21], and have been extended to multicast scenarios [22], as well as to two-timescale designs exploiting statistical channel state information (CSI) [23]. These and other related works on MA-enabled systems demonstrate that adding mechanical DoF to electronic beamforming can significantly improve performance [24, 25, 26, 27].
@@
However, most existing MA- or RIS-related works treat antenna movement or passive phase tuning in isolation [28], and their ISAC designs for LAWN-like geometries are much less explored. Existing ISAC designs typically assume fixed arrays and do not exploit low-altitude 3D geometry control to shape the transmit beampattern and the multiuser downlink channels simultaneously [29]. Furthermore, the mechanical flexibility in the MA literature is typically modeled as small-scale position changes within a local region. In contrast, realistic low-altitude platforms, e.g., UAV-mounted arrays or rooftop BS panels, can be designed to support rotation of the entire active or passive array via low-cost mechanical actuators [30, 31, 32]. Such rotatable arrays (RA) can jointly adjust their boresight directions, element coordinates, and directional gains, offering a distinct and practically relevant form of geometry control [33, 34]. In summary, RIS-aided ISAC designs that consider rotatable passive arrays and their coupling with active-array rotations at the BS have yet to be systematically studied, especially under explicit constraints on the similarity of sensing beampatterns. As a result, the fundamental question of how to jointly design rotatable active and passive arrays, RIS phase shifts, and multiuser transmit beamforming for low-altitude ISAC remains open [35].
@@
Motivated by these observations, this paper investigates a low-altitude ISAC architecture with cooperatively rotatable active and passive arrays. We consider a downlink LAWN scenario in which a BS equipped with an active rotatable array simultaneously serves multiple single-antenna ground users and senses low-altitude targets in monostatic mode, assisted by a rotatable RIS, termed a passive array. Both arrays can rotate in 3D, and the resulting orientations jointly affect the element coordinates and equivalent phases in the steering vectors and direction-dependent array gains through a practical radiation model. Based on the rotation-aware channel model, we formulate a joint design problem on the BS precoder, the RIS phase shift vector, and the rotation angles of arrays, which maximizes a weighted-sum ISAC utility of the downlink sum rate and the sensing beampattern similarity, subject to the BS transmit-power budget, RIS unit-modulus constraints, and mechanical rotation limits at both the BS and the RIS. The main contributions of this work are summarized as follows
@@
  • We propose a LAWN-oriented ISAC architecture where a rotatable active BS array cooperates with a rotatable passive RIS array. We develop a rotation-aware 3D multipath channel model that captures both phase shifts of steering vectors induced from rotated antenna and element positions and directional path gains induced from direction-dependent antenna and element radiation patterns. We formulate an optimization problem that maximizes the weighted-sum utility on communication sum rate and sensing beampattern similarity, which is characterized by normalized mean-squared error (NMSE) between the generated and desired beampattern, by jointly designing BS transmit beamforming, RIS phase shifts, and active and passive array rotations, subjected to power budget on the BS, unit-modulus constraint on the RIS, and mechanical rotation bounds on both RAs.
@@
  • We develop an alternating optimization (AO) framework that iteratively updates the BS precoder, RIS phase shifts, and array rotations. For the BS precoding block, we combine the quadratic transform (QT) and the Lagrangian dual transform (LDT) with a majorization–minimization (MM) step to obtain a tractable quadratic-constrained quadratic program (QCQP) subproblem that admits a closed-form water-filling-like update. For the RIS block, we exploit the complex-circle manifold structure and design an efficient Riemannian conjugate gradient (RCG) algorithm. For the rotation block, we derive closed-form gradients of both SINR and beampattern NMSE with respect to the BS and RIS Euler angles and employ a projected gradient descent (PGD) method with Barzilai–Borwein initialization and Armijo backtracking.
@@
  • We conduct extensive simulations using LAWN-inspired 3D geometries and compare the proposed design with several baselines, including only BS rotation, only RIS rotation, fully fixed arrays, and non-RIS configurations. The results show that the jointly rotatable BS and RIS architecture provides substantial gains in the weighted ISAC utility of the communication rate and beampattern NMSE across varying transmit powers, antenna array sizes, and user densities. The gains are further amplified by practical directional radiation patterns, highlighting the importance of aligning the array boresights and main lobes with both the user and target directions in low-altitude 3D space. By sweeping the trade-off weight, we further characterize the Pareto frontier between the sum rate and the NMSE, thereby quantifying how joint array rotations expand the ISAC trade-off boundary.
@@
The rest of the paper is structured as follows: Section II introduces the system model and problem formulation. Section III presents the proposed algorithm. Section IV presents the numerical results, and Section V concludes the paper.
Notations: Scalars are denoted by italic letters, and vectors and matrices are denoted by bold-face lower and upper-case letters, respectively. $\mathbb{R}^{x}$ and $\mathbb{R}^{x\times y}$ denote the space of real vectors or matrices, and $\mathbb{C}^{x}$ and $\mathbb{C}^{x\times y}$ denote the space of complex vectors or matrices, respectively. $\left|x\right|$ denotes the modulus of a complex-valued scalar $x$. For a vector $\boldsymbol{x}$, $\left\|\boldsymbol{x}\right\|$ denotes its Euclidean norm and $\boldsymbol{x}^{*}$ denotes its conjugate. The distribution of a circularly symmetric complex Gaussian random vector with mean vector $\boldsymbol{x}$ and covariance matrix $\boldsymbol{\varSigma}$ is denoted by $\mathcal{CN}\left(\boldsymbol{x},\boldsymbol{\varSigma}\right)$. The Euclidean gradient of a scalar function $f(\boldsymbol{x})$ with a vector variable $\boldsymbol{x}$ is denoted by $\nabla f(\boldsymbol{x})$. For a square matrix $\boldsymbol{S}$, $\mathrm{Tr}\left(\boldsymbol{S}\right)$ denote its trace. For any general matrix, $\boldsymbol{M}^{H}$ and $\left[\boldsymbol{M}\right]_{ij}$ denote its conjugate transpose and $(i,j)$th element,respectively. $\boldsymbol{I}$ denotes an identity matrix, respectively. $\mathbb{E}[\cdot]$ stands for the expectation operation.
@@
Figure 1: A low-altitude ISAC system with rotatable active and passive arrays.
@@ II System Model and Problem Formulation
We consider a downlink RA-enabled ISAC system, as shown in Fig. 1, where a BS uses an active rotatable uniform planar array (UPA) with $M$ transmit antennas to serve $K$ single-antenna users while sensing multiple targets in monostatic mode; at the same time, a passive rotatable RIS arranged as a UPA with $N$ reflecting elements assists both tasks. In addition to active transmit beamforming at the BS and passive phase shifts at the RIS, both arrays can rotate in 3D. These rotations steer the array boresights and reshape the array response to improve or balance the performance between communication and sensing.
@@ II-A Array Geometry and Rotations
We adopt a global right-handed Cartesian coordinate system (CCS) $\mathcal{G}=(O,\mathrm{x},\mathrm{y},\mathrm{z})$. The BS employs a co-located transmit/receive UPA with size $M=M_{\text{col}}\times M_{\text{row}}$ of $M_{\text{col}}$ columns and $M_{\text{row}}$ rows, and the RIS is also a UPA with size $N=N_{\text{col}}\times N_{\text{row}}$ of $N_{\text{col}}$ columns and $N_{\text{row}}$ rows. Let $\overline{\boldsymbol{d}}^{\text{B}}_{m}\in\mathbb{R}^{3}$ and $\overline{\boldsymbol{d}}^{\text{R}}_{n}\in\mathbb{R}^{3}$ denote the unrotated local coordinates of the $m$-th BS transmit antenna and the $n$-th RIS reflecting element, respectively, and let $\boldsymbol{d}_{0}^{\text{B}}$ and $\boldsymbol{d}_{0}^{\text{R}}$ be their array centers in $\mathcal{G}$. Both BS and RIS employ a motorized mechanism to rotate the arrays as a whole. Based on the system’s ISAC requirements, the proper rotation angles are determined by a central control center, which is connected to the motor controllers via fiber links [^1] In practice, mechanical rotations can be updated on a slow timescale (e.g., sector-level adaptation or quasi-static geometry changes), whereas the BS digital precoder and RIS phases can be refreshed on the fast timescale of each channel-coherence block. Since the mechanical control involves only six scalar angles (3D Euler angles for the BS and RIS), the control signaling overhead is negligible compared with per-element electronic tuning. When the angles change, the array responses are updated deterministically; thus, standard pilot-based estimation and tracking remain applicable, and two-timescale designs using long-term statistical CSI for rotations are also feasible.. Both the active array at BS and the passive RIS array rotate as a rigid body through Euler angles denoted by
$\boldsymbol{r}^{\text{B}}\triangleq\big[r_{\mathrm{x}}^{\text{B}},r_{\mathrm{y}}^{\text{B}},r_{\mathrm{z}}^{\text{B}}\big]^{T},\quad\boldsymbol{r}^{\text{R}}\triangleq\big[r_{\mathrm{x}}^{\text{R}},r_{\mathrm{y}}^{\text{R}},r_{\mathrm{z}}^{\text{R}}\big]^{T},$ (1)
respectively, in their local CCS. In the following, we denote $\text{A}\in\{\text{B, R}\}$ as the notations for either BS or RIS array for brevity. Then, the corresponding rotation matrix generated by $\boldsymbol{r}^{\text{A}}$ represents the rotations about their local axis with the standard intrinsic $\mathrm{x}$-$\mathrm{y}$-$\mathrm{z}$ convention and is given by
$\boldsymbol{R}\left(r_{\mathrm{x}}^{\text{A}},r_{\mathrm{y}}^{\text{A}},r_{\mathrm{z}}^{\text{A}}\right)=\boldsymbol{R}_{\mathrm{x}}\left(r_{\mathrm{x}}^{\text{A}}\right)\boldsymbol{R}_{\mathrm{y}}\left(r_{\mathrm{y}}^{\text{A}}\right)\boldsymbol{R}_{\mathrm{z}}\left(r_{\mathrm{z}}^{\text{A}}\right),$ (2)
@@
$\boldsymbol{R}_{\mathrm{x}}\left(r_{\mathrm{x}}^{\text{A}}\right)=\left[\begin{matrix}1&0&0\\ 0&\cos r_{\mathrm{x}}^{\text{A}}&-\sin r_{\mathrm{x}}^{\text{A}}\\ 0&\sin r_{\mathrm{x}}^{\text{A}}&\cos r_{\mathrm{x}}^{\text{A}}\\ \end{matrix}\right],$ (3a)
$\boldsymbol{R}_{\mathrm{y}}\left(r_{\mathrm{y}}^{\text{A}}\right)=\left[\begin{matrix}\cos r_{\mathrm{y}}^{\text{A}}&0&\sin r_{\mathrm{y}}^{\text{A}}\\ 0&1&0\\ -\sin r_{\mathrm{y}}^{\text{A}}&0&\cos r_{\mathrm{y}}^{\text{A}}\\ \end{matrix}\right],$ (3b)
$\boldsymbol{R}_{\mathrm{z}}\left(r_{\mathrm{z}}^{\text{A}}\right)=\left[\begin{matrix}\cos r_{\mathrm{z}}^{\text{A}}&-\sin r_{\mathrm{z}}^{\text{A}}&0\\ \sin r_{\mathrm{z}}^{\text{A}}&\cos r_{\mathrm{z}}^{\text{A}}&0\\ 0&0&1\\ \end{matrix}\right].$ (3c)
The rotated antenna or element coordinates in $\mathcal{G}$ are then